OpenStudy (anonymous):

A store sold towels that cost $5 each at 15% less than the cost. The overhead was 20% of the cost of each towel. WHat was the selling price of each towel? How much more than $150 did the owner lose if 95 toels were sold?

OpenStudy (anonymous):

May someone please help me!

hero (hero):

I don't know what overhead means.

OpenStudy (anonymous):

Overhead- includes things as advertising, rent, utilities

hero (hero):

The selling price shouldn't have anything to do with the overhead

OpenStudy (anonymous):

Well that's the question

hero (hero):

Okay, think I get it now.

OpenStudy (anonymous):

So he's selling it at 15% less than what he bought it for (why he would do that I have no idea):\[selling price = (1-0.15) \times 5.00 = 0.85 \times 5.00 = 4.25\] So that's you're selling price ($4.25). The overhead (meaning money spent/lost per item by the owner) is 20% of the cost:\[overhead = 0.2 \times 5.00 = 1\] So for every towel, he's also losing $1. When you include the loss from selling the towels below cost, he's losing $1.75 per towel. So to find out how much he lost by selling 95 towels:\[loss = 1.75 \times 95 = 166.25\] They ask how much above $150 he loses, so that would be\[166.25 - 150.00 = 16.25\]

OpenStudy (anonymous):

Also, you should click the "Good Answer" button by my name if my answer helps you.